EA012B

PI: Alef
Title: Andromedae

Session: Oct 1996 (session 3)
Frequency: 3.6 cm
Mode: A
Obs time (UT): 304 EU 14:00(30/10)-04:00(31/10)
Stations: EbNtOnShUr
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (J. Conway, A. Polatidis at Mon Nov 4 11:14:09 EST 1996) Success
No known problems. Sky mostly clear during the longer part of the experiment.

Shanghai (Huang Xinyong at Tue Nov 5 03:53:41 EST 1996) Success
No problems. Clock offset 0.1 micro sec. Weather was ovsercast.

Onsala (J. Conway, A. Polatidis at Tue Nov 5 05:32:57 EST 1996) Success
Problems with placing of the READ head between UT 163538-163739 (Day 304). Data should be OK. No other known problems.

OAN\-Yebes (Pablo de Vicente at Tue Nov 5 08:53:07 EST 1996) Minor failures
Scans from 02:14:10 - 03:38:26 lost due to the antenna control computer

Noto (Corrado Trigilio at Tue Nov 5 11:42:04 EST 1996) Success
No problems reported. Clear sky.

Effelsberg (d graham at Wed Nov 6 09:25:38 EST 1996) Minor failures
1907 to 2122 no data, telescope would not drive

Urumqi (Liu Xiang at Tue Nov 19 03:46:16 EST 1996) Minor failures
UT14:00:00-14:37:56, was not recorded since head position failed. Clock offset 0.2+/-0.4us.

MPIfR (Walter Alef at Tue Sep 23 17:29:05 METDST 1997) Minor failures
Urumqi gave no fringes. Last hour of Crimea lost. Rest O.K. Fringes for: Eb, Sh, Sm, Nt, On, Ye, Wz, Ro, Sm.